OpenStudy (anonymous):

Write the equation of the hyperbola. Foci: (-3,0), (1,0) Vertices: (-2,0), (0,0)

OpenStudy (anonymous):

yikes these are endless

OpenStudy (anonymous):

the center is half way between the vertices, so it must be \((-1,0)\)

OpenStudy (anonymous):

that means it is going to look like \[\frac{(x+1)^2}{a^2}-\frac{y^2}{b^2}=1\] now we need \(a\) and \(b\)

OpenStudy (anonymous):

the vertex is one unit from the center, so \(a=1\) and of course \(a^2=1\)

OpenStudy (anonymous):

the focus is 2 units from the center so \(c=2\) and \(b^2=2^1-1^2=3\)

OpenStudy (anonymous):

final answer is \[(x+1)^2-\frac{y^2}{3}=1\]

OpenStudy (anonymous):

wow thank you so much, you have no idea how much this helps me

OpenStudy (anonymous):

yw, again all you really need to do is look here http://www.purplemath.com/modules/hyperbola3.htm and follow along making the necessary changes with the numbers

OpenStudy (anonymous):

ill try that :)

OpenStudy (anonymous):

you can check the answer here http://www.wolframalpha.com/input/?i=hyperbola+%28x%2B1%29^2-y^2%2F3%3D1